OpenStudy (anonymous):

given V=-3000t+21,000 can be used to find the value V of a car at the end of t years? In how many years is the car worth 12,000?

OpenStudy (mathteacher1729):

The value of the car after t years is \[V(t)=21,000 - 3000t\] In other words, the car starts off being worth $21,000 and it loses $3,000 in value every year. If you want to know how many years it takes for the car to reach $12,000, then solve: \[12000 = 21,000 - 3000t\] You can do this in your head, because everything is divisible by 1,000. :) \[12 = 21 - 3t\] Now just solve for t. :) Hope this helps!

OpenStudy (anonymous):

you would multiple t by 3000 and minuse it by the original sum(21000)

OpenStudy (anonymous):

an example is if t is 2 then the forumal would look like this V=-3000x2+21,000 v=15000

OpenStudy (anonymous):

the year was given previously as 5 years.

OpenStudy (anonymous):

12=21-3(11) would get a negative 12 but i need it to be postive

OpenStudy (anonymous):

if it was5 years then you would use the original forumila and instead of t you would put 5

OpenStudy (anonymous):

yes that would give it 15 years to be the answer

OpenStudy (anonymous):

do u understand it now or do u need some more help

OpenStudy (mathteacher1729):

5, 15, and 11 years are NOT the answers. If you subsitute those into the original equation given, they do not give $12,000 as the price of the car.

OpenStudy (anonymous):

21-3(3)=12

OpenStudy (anonymous):

I needed to figure out if its true the car will be worth 200 in 7 years would i make the equation to be 200=-3000(7)+21000?